A given mass of a gas occupies 850cm3 at 320K and 0.92×105 Nm-2 pressure. Calculate the volume of the gas at S.T.P.
1
Expert's answer
2020-07-07T14:35:21-0400

P1 = 0.92×105 N m-2 = 0.92×105 Pa = 92000 Pa

V1 = 850 cm3 = 0.85 L

T1 = 320 K

P2 = 101325 Pa (STP)

V2 = unknown

T2 = 273.15 K (STP)

 

Solution:

The Combined Gas Law:

P1V1/T1 = P2V2/T2

Cross-multiply to clear the fractions:

P1V1T2 = P2V2T1

Divide to isolate V2:

V2 = (P1V1T2) / (P2T1)

Plug in the numbers and solve for V2:

V2 = (92000 Pa × 0.85 L × 273.15 K) / (101325 Pa × 320 K) = 0.659 L = 659 mL = 659 cm3

V2 = 659 cm3


Answer: 659 cm3 is the volume of the gas at STP.